OpenStudy (anonymous):

Eric eats 1/6 of an apple pie and Jack eats 1/4 of the appel pie. How much of the pie has been eaten? A 1/5 B 5/12 C 7/12 D 1/3

OpenStudy (karabella98):

You have to add the fractions, but to do that you have to find the least common denominator

OpenStudy (anonymous):

The equation would look like this \[\frac{ 1 }{ 6 } + \frac{ 1 }{ 4 }\] However, you need to find the LCD before you add.

OpenStudy (anonymous):

it would be 5/12 cause the lease common deminator is 12 cause 6*2 and 4*3

OpenStudy (anonymous):

The LCD would be 12. So, multiply 1 and 6 by 2, and multiply 4 and 1 by 3. You'd get: \[\frac{ 2 }{ 12 } + \frac{ 3 }{ 12 }\] Now just add the fractions. 2 + 3 = 5 So, the answer is 5/12.